OpenStudy (anonymous):

Help in Algebra 1? Question and choises will be in comments

OpenStudy (anonymous):

Which expression is the simplified form of \[\sqrt{150b^2}\] A) \[6b \sqrt{5b}\] B)\[5b \sqrt{\sqrt{6}}\] C)\[6b \sqrt{5}\] D)\[5b \sqrt{6b}\]

OpenStudy (jim766):

\[5b \sqrt{6}\]

OpenStudy (anonymous):

Thank You!